OpenStudy (anonymous):

standard deviation of 2 4 8 12 9

OpenStudy (anonymous):

Step one: Find the average of the data. So, add all the numbers up, then divide by how many numbers are there. So, \[\frac{ 2+4+8+12+9 }{ 5 } = 7\]

OpenStudy (anonymous):

Step Two: Take the average, which is 7, then subtract it from all of the numbers there. So: \[2-7 = -5\] \[4-7=-3\] \[8-7=1\] \[12-7=5\] \[9-7=2\]

OpenStudy (anonymous):

Step Three: Square each value we just found. \[-5^{2}=25\] \[-3^{2}= 9\] \[1^{2}=1\] \[5^{2}=25\] \[2^{2}=4\]

OpenStudy (anonymous):

Step four: Add all of those together. \[25+9+1+25+4=64\]

OpenStudy (anonymous):

Step Five; Then divide that number by the amount of numbers in the data, and take away one. So, the amount of numbers there is 5. take one away we have 4. Then we do 64/4. \[\frac{ 64 }{ 4 } = 16\]

OpenStudy (anonymous):

Last Step. We take the square root. \[\sqrt{16}= 4\]

OpenStudy (anonymous):

The final answer is 4!

OpenStudy (anonymous):

Do you understand?